# Towards Robust Semantic Segmentation against Patch-based Attack via Attention Refinement

### v1

@@ 1 Introduction
1. We are the first to analyze the robustness of the attention-based semantic segmentation model against patch-based attacks.
2. Through the analysis of the effective receptive field, we attribute the vulnerability of models against patch-based attacks to the attention mechanism used in semantic segmentation models.
@@
Figure 1: A visualization example of the attack setting of Permute and Strip. Permute refers to the setting that replaces the labels in the ground truth segmentation with labels that do not appear in them one by one. Strip refers to the setting that chooses a strip pattern consisting of randomly selected labels as the target segmentation label.
@@ 2.1 Semantic Segmentation
Early semantic segmentation methods use CNN models as backbones to extract features. FCN [27] is the first to propose fully convolutional networks, which can adapt to any resolution of input images. Since CNN extracts features by stacking multiple layers of convolution, there is a contradiction between preserving the receptive field and preserving resolution in earlier works. DeepLab [5] proposes a dilated convolution layer. It enlarges the receptive field of networks without introducing additional parameters. And some works [6, 28] adopt multi-scale feature aggregation modules to guarantee both the large-scale receptive field and high-resolution feature map. Self-attention [29] is another important module that introduces global receptive filed for each pixel, and its subsequent works [30, 31, 32] adopt various methods to reduce the complexity. Based on self-attention, vision transformer [26] has become the mainstream model in various computer vision tasks, and its variants in semantic segmentation [33, 25, 34, 23, 35, 36] also achieve promising results.
@@ 3.2 Analysis
The visualization of the effective receptive field of different models and the corresponding robustness against patch-based attacks are shown in Table 1. In the experiments, we evaluate the robustness of models against two different attack targets, named Permute and Strip, respectively. A visualization example of the two attack settings is shown in Fig. 1 and the details of attack settings are described in Section 5.1. For the patch-based attack, we use the method of DAG [18] by adding an adversarial patch with the size of $150\times 150$ in the lower right corner of the image. We use the mIoU between the segmentation results of the adversarial examples and the target labels determined by the adversary as the metric to evaluate the robustness of the semantic segmentation model against these attacks. Lower mIoU values indicate the better robustness of the model.
@@
It is worth mentioning that Mao et al. [67] find that compared with the global attention mechanism in ViT [26], the local attention used in Swin [24] damages the robustness of classification models against pixel-wise attacks. It is opposite to the conclusion we draw for the robustness of semantic segmentation models against patch-based attacks in our work. It reflects the unique value of the robustness research on semantic segmentation models under the patch-based attack settings, and highlights the contributions of our work.
@@ 5.1 Experimental Setting
Attack settings. In the experiments, we set up two different kinds of attack targets, named Permute and Strip, respectively. Suppose the ground truth segmentation label corresponding to the input image $\mathbf{X}$ is $\mathbf{Y^{g}}\in\{0,1,\cdots,C-1\}^{H\times W}$ and the target segmentation label of the attack is $\mathbf{Y^{t}}\in\{0,1,\cdots,C-1\}^{H\times W}$. Suppose the set of categories that appears in $\mathbf{Y^{g}}$ is $S=\{y_{1},...,y_{n}\}\subseteq\{0,1,\cdots,C-1\}$.
@@
A visualization example of the two attack settings is shown in Fig. 1.
Metrics. Following previous semantic segmentation works, we adopt mean Intersection-over-Union (mIoU) as the metric to evaluate the semantic segmentation model for clean images. Besides, to assess the robustness of models under two attack settings, we compare the segmentation results of the adversarial examples and the target labels determined by the adversary, and calculate their mIoU as the metric. Here, the smaller the mIoU means the robustness of the model is better. In order to fully demonstrate the effectiveness of our method, we also use the metric of pixel accuracy (pAcc) to evaluate the robustness of different models. Specifically, we calculate pAcc as follows:
@@ 5.2 Ablation Study
MAS RAD (mIoU/%) (mIoU/%)
$\times$ $\times$ 49.32 40.01
$\times$ $\surd$ 46.77 37.83
$\surd$ $\times$ 34.11 23.36
$\surd$ $\surd$ 33.72 22.69
@@ 5.3 The Impact of the Hyperparameter
In this section, we conduct the experiments to confirm the optimal hyperparameters in our proposed RAM, e.g., the threshold $T$ in Max Attention Suppression and the dropout rate $p$ in Random Attention Dropout. It can be seen from Table 4 and Table 5 that as the threshold $T$ and dropout rate $p$ becomes larger, the robustness of the model first becomes better and then worse. When the threshold and dropout rate are 0.3 and 0.5 respectively, the robustness of the model against patch-based attack is the best. In all other experiments, we set the threshold $T$ in MAS to 0.3 and the dropout rate $p$ in RAD to 0.5.
@@ 5.6 Comparison with Existing Methods
From Table 8, we can also find that although the clean mIoU of the model may be slightly reduced after combined with RAM, the sacrifice is acceptable and negligible considering the huge improvement of the model robustness brought by the RAM method.
@@ 5.8 Comparison with Adversarial Training
We also compare our proposed defense method with the method of adversarial training. We use the method of PGD [13] with 10 attack steps to conduct the adversarial training. It should be noted that, although adversarial training is a widely used and effective defense method, the training time and computing resources required by adversarial training far exceed our method of refining the attention mechanism. From Table 11, it can be seen from the table that under almost all attack methods, our method achieves better robustness than the adversarial training method, which further demonstrates the effectiveness of our proposed defense method.
@@ 5.10 Visualization
We visualize the segmentation results of Nonlocal [29]/R50 [1] model with our proposed RAM and the baseline model under the attack setting of Permute. As shown in Fig. 3, most of the segmentation results towards the adversarial examples of our proposed RAM model still maintain the ground truth label. In contrast, the adversarial example almost completely modifies the output of the baseline model, and the segmentation results are very close to the target labels specified by the adversary. It demonstrates that the baseline model is quite vulnerable to patch-based attacks, while our RAM method significantly improves the robustness of semantic segmentation models against attacks. More visualization results are shown in Appendix B.
@@ A.1 The General Framework of Attack Methods
In the experiments, we set $T=400$ for all attack methods. For the hyperparameter of $\gamma$, we set $\gamma=0.005$ for PGD [13], and $\gamma=1$ for the other three attack methods.

### v2

@@ 1 Introduction
1. Through the analysis of the effective receptive field, we conclude that the vulnerability of models against patch-based attacks is attributed to the attention mechanism used in semantic segmentation models.
@@
Figure 1: A visualization example of the targeted attack setting of Permute and Strip. Permute refers to the setting that replaces the labels in the ground truth segmentation with labels that do not appear in them one by one. Strip refers to the setting that chooses a strip pattern consisting of randomly selected labels as the target segmentation label.
@@ 2.1 Semantic Segmentation
Early semantic segmentation methods use CNN models as backbones to extract features. FCN [27] is the first to propose fully convolutional networks, which can adapt to any resolution of input images. Since CNN extracts features by stacking multiple layers of convolution, there is a contradiction between preserving the receptive field and preserving resolution in earlier works. DeepLab [5] proposes a dilated convolution layer. It enlarges the receptive field of networks without introducing additional parameters. And some works [6, 28] adopt multi-scale feature aggregation modules to guarantee both the large-scale receptive field and high-resolution feature map. Self-attention [29] is another important module that introduces global receptive filed for each pixel, and its subsequent works [30, 31, 32] adopt various methods to reduce the complexity. Based on self-attention, vision transformer [26] has become the mainstream model in various computer vision tasks, and its variants in semantic segmentation [33, 25, 34, 23, 35, 36] also achieve promising results. Recently, SegViT v2 [37] introduces a novel Attention-to-Mask (ATM) module to design a lightweight decoder effective for plain ViT.
@@ 3.2 Analysis
The visualization of the effective receptive field of different models and the corresponding robustness against patch-based attacks are shown in Table 1. In the experiments, we evaluate the robustness of models against two different attack targets, named Permute and Strip, respectively. A visualization example of the two targeted attack settings is shown in Fig. 1 and the details of attack settings are described in Section 5.1. For the patch-based attack, we use the method of DAG [18] by adding an adversarial patch with the size of $150\times 150$ in the lower right corner of the image. We use the mIoU between the segmentation results of the adversarial examples and the target labels determined by the adversary as the metric to evaluate the robustness of the semantic segmentation model against these attacks. Lower mIoU values indicate the better robustness of the model.
@@
It is worth mentioning that Mao et al. [68] find that compared with the global attention mechanism in ViT [26], the local attention used in Swin [24] damages the robustness of classification models against pixel-wise attacks. It is opposite to the conclusion we draw for the robustness of semantic segmentation models against patch-based attacks in our work. It reflects the unique value of the robustness research on semantic segmentation models under the patch-based attack settings, and highlights the contributions of our work. We attribute the opposite conclusions to the combined effects of different attack types (pixel-wise vs. patch-based) and different task models (classification vs. segmentation). In classification tasks, under pixel-wise attacks, the entire image is subjected to noise interference. To make a correct classification, models may not need the information from the entire image, but can rely on distinctive local content for decision-making. When the critical discriminative region is disrupted, global attention methods, with their larger receptive fields, can still generate the correct label by relying on information from other locations. For segmentation tasks during patch-based attacks, the scenario differs as segmentation requires classifying all pixels in an image. When an image suffers a patch-based attack, the adversarial perturbation appears only in a specific part of the image. Global attention might spread the localized disturbance across the entire image, potentially affecting the accuracy of segmentation results in more areas.
@@ 4.3 Random Attention Dropout
4.4 Discussion
In this section, we will discuss in detail the differences between our work and the related study [69]. Firstly, our work and [69] focus robustness analysis on different tasks: [69] investi-gates the robustness of attention mechanisms in classification models against patch-based perturbations. Our research, however, studies the robustness of attention mechanisms in semantic segmentation models to patch-based perturbations. Given the distinct tasks of segmentation and classification models, patch-based attacks on segmentation tasks have unique characteristics. Semantic segmentation models are designed to provide segmen-tation results for every pixel within an entire image, where a local dirty patch could sig-nificantly influence the segmentation results through the attention mechanism, spreading its negative effect throughout the image. In contrast, image classification results may rely on features from key areas. Therefore, in image classification tasks, local dirty patches may not have a particularly significant impact on the overall classification result. Also, our work differs from [69] in methodology: [69] suggests replacing the original attention with smoothed attention via temperature scaling in the softmax operation. Conversely, our method reduces the potential impact of dirty patches on other areas by compressing the value range of the attention map through max attention suppression and mitigates the negative effects of dirty patches by randomly dropping some patches using random attention dropout.
@@ 5.1 Experimental Setting
Attack settings. In the experiments, we conduct attacks under both targeted and untargered settings. For the targeted attack setting, we set up two different kinds of attack targets, named Permute and Strip, respectively. Suppose the ground truth segmentation label corresponding to the input image $\mathbf{X}$ is $\mathbf{Y^{g}}\in\{0,1,\cdots,C-1\}^{H\times W}$ and the target segmentation label of the attack is $\mathbf{Y^{t}}\in\{0,1,\cdots,C-1\}^{H\times W}$. Suppose the set of categories that appears in $\mathbf{Y^{g}}$ is $S=\{y_{1},...,y_{n}\}\subseteq\{0,1,\cdots,C-1\}$.
@@
A visualization example of the two targeted attack settings is shown in Fig. 1. It should be noted that our reason for adopting targeted attack settings, in addition to the common untargeted attack scenarios, is to simulate more practical scenarios. In the context of untargeted attacks, the segmentation results of the perturbed images are significantly disrupted, making it easier for the model providers to visually inspect the model outputs and detect the model is under attack, thereby enabling the enactment of specific countermeasures. Consequently, we resort to the targeted attack and use target labels designated by the attackers themselves, producing adversarial examples (e.g., under the Permute setting) that, although bearing incorrect semantic category labels, remain shape-wise similar to the ground truth labels (as shown in Fig. 1), i.e., the high similarity between edge contours of ground-truth and the adversarial labels, making them relatively harder for model providers to visually detect.
Metrics. Following previous semantic segmentation works, we adopt mean Intersection-over-Union (mIoU) as the metric to evaluate the semantic segmentation model for clean images. Besides, to assess the robustness of models under two targeted attack settings, we compare the segmentation results of the adversarial examples and the target labels determined by the adversary, and calculate their mIoU as the metric. Here, the smaller the mIoU means the robustness of the model is better. For the untargeted attack setting, we use the metric of the mIoU between the segmentation results of the adversarial images and the ground truth labels to evaluate the robustness of the model. In order to fully demonstrate the effectiveness of our method, we also use the metric of pixel accuracy (pAcc) to evaluate the robustness of different models. Specifically, we calculate pAcc as follows:
@@ 5.2 Ablation Study
MAS RAD (mIoU/%) (mIoU/%) (mIoU/%)
$\times$ $\times$ 48.25 49.32 40.01
$\times$ $\surd$ 48.37 46.77 37.83
$\surd$ $\times$ 48.10 34.11 23.36
$\surd$ $\surd$ 48.30 33.72 22.63
@@ 5.3 The Impact of the Hyperparameter
In this section, we conduct the experiments to confirm the optimal hyperparameters in our proposed RAM, e.g., the threshold $T$ in Max Attention Suppression and the dropout rate $p$ in Random Attention Dropout. As $T$ decreases, the upper limit of the attention value associated with the dirty patch decreases, consequently mitigating its influence on other patches. However, when $T$ is too small, the compressed range of the attention map values $[0,T]$ becomes too limited, leading to a lack of differentiation between features of different patches, which could negatively affect the model’s performance. Therefore, as shown in Table 4, as $T$ decreases, the results of our method initially improve but then deteriorate. Through ablation experiments, we have ultimately chosen $T=0.3$ as the hyperparameter. The parameter $p$ is the dropout rate if our proposed RAD method. The larger $p$ is, the greater the number of patches randomly dropped, increasing the probability of the dirty patch being randomly discarded, thereby reducing its impact on other clean patch locations. On the other hand, discarding more clean patches can affect the segmentation results. Hence, as shown in Table 5, as $p$ increases, the robustness of the model initially improves but then decreases. Ultimately, we select $p=0.5$ as the hyperparameter.
@@ 5.6.1 Targeted Attack Setting
From Table 8, we can also find that although the clean mIoU of the model may be slightly reduced after combined with RAM. Our proposed MAS module can reduce the impact of dirty patches on the features of other locations by compressing the value range in the attention map. However, when dealing with clean images, where no dirty patches are present, the compression of the value range of the attention map within MAS can diminish the distinction between the features of different patches, which may negatively affect the clean performance of model. Yet, our method significantly enhances the model’s robustness to patch-based attacks. Considering the substantial improvement in robustness, the decrease in clean accuracy is considered negligible.
@@
5.6.2 Untargeted Attack Setting
We also conduct experiments to compare the robustness of different models under the untargeted attack setting. We utilize adversarial patches of size $100\times 100$, positioned at the bottom right corner of the images. Our experiments are conducted on the Cityscapes dataset [73] using the DAG attack method [18] to generate adversarial examples. We use the metric of the mIoU between the segmentation results of the adversarial images and the ground truth labels. The higher mIoU means the better robustness of the model. The results are shown in Table 10. From the results presented in the table, it’s observed that our proposed method achieves an improvement of over 10% in mIoU compared to the baseline across different models. Compared to existing defense methods, our approach also demonstrates a significant improvement in robustness, verifying the effectiveness of our method again.
Table 10: The robustness of different semantic segmentation models against patch-based untargeted attack settings. clean represents the mIoU of the semantic segmentation results on the validation set, and attack represent the mIoU between the segmentation results of the adversarial images and the ground truth labels of the clean images. The higher mIoU regards to clean means the higher segmentation accuracy of the model for clean images, and the higher mIoU regards to attack means the better robustness of the model. The attack is conducted by DAG [18].
Model Global Attention Local Attention
$\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
(mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%)
Baseline 76.12 25.49 78.56 23.62 79.10 37.81 78.91 35.62
Learnable [68] 75.37 24.63 77.82 25.47 77.32 38.74 78.01 36.32
Temperature [69] 75.42 27.52 77.93 24.92 77.19 39.93 77.42 39.17
RAM (ours) 75.91 35.56 78.03 32.67 78.09 50.97 77.78 48.31
@@ 5.8 Comparison with Adversarial Training
We compare our proposed defense method with the method of adversarial training under two targeted attack settings. We use the method of PGD [13] with 10 attack steps to conduct the adversarial training. It should be noted that, although adversarial training is a widely used and effective defense method, the training time and computing resources required by adversarial training far exceed our method of refining the attention mechanism. From Table 12, it can be seen from the table that under almost all attack methods, our method achieves better robustness than the adversarial training method, which further demonstrates the effectiveness of our proposed defense method.
@@
We also compare our RAM with advanced adversarial training method [75] under the untargeted attack setting in Table 13. From the table we can find that the robustness of the DDC-AT method [75] is similar to our proposed RAM, but it is important to note that DDC-AT require approximately three times the training time compared to our method. Additionally, our proposed method can be seamlessly tailored into adversarial training approaches to further boost the robustness of semantic segmentation models against patch-based attacks. As seen in the table, combining our RAM with DDC-AT further improves the robustness of semantic segmentation models to patch-based attacks.
Table 13: Integration with advanced adversarial training method.
Model Global Attention Local Attention
$\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
(mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%) (mIoU/%)
Baseline 76.12 25.49 78.56 23.62 79.10 37.81 78.91 35.62
DDC-AT [75] 75.21 34.78 76.86 32.88 77.10 49.69 76.28 48.16
RAM (ours) 75.91 35.56 78.03 32.67 78.09 50.97 77.78 48.31
DDC-AT [75] + RAM (ours) 75.37 37.81 77.74 33.52 77.65 52.43 77.46 50.97
@@ 5.10 Visualization
We compare the visualizations of attention maps when facing adversarial examples between our proposed RAM method and baseline methods, as shown in Fig. 3. We conduct experiments on the ADE20K dataset. We compare the attention maps of the feature at the adversarial patch location to all locations in the image, generated using different methods. In the figure, the first column shows the clean image, the second column shows the attention map of the clean image, the third and fourth columns show the attention maps of the feature at the adversarial patch location to all locations in the image, generated by the baseline method and our proposed RAM method, respectively. After converting the attention map into a heatmap, we overlay it on the corresponding adversarial example with an alpha of 0.3 for visualization purposes. Redder colors in the foreground indicate greater influence. From the figure, it is evident that in the clean image, local patches that are semantically relevant to specific regions exhibit greater attention, whereas regions that are semantically irrelevant display lesser attention. In contrast, the influence of the adversarial patch spreads to all areas, regardless of their semantic relevance to the region’s content. Additionally, we can observe that our proposed RAD module, by suppressing the value range of the attention map, significantly reduces the impact of adversarial patch features on the features of other clean areas compared to baseline methods. Therefore, it can enhance the robustness of semantic segmentation models to patch-based attacks.
Figure 3: Visualization of attention maps, where the first column shows the clean image, the second column shows the attention map of the clean image, the third and fourth columns show the attention maps of the feature at the adversarial patch location to all locations in the image, generated by the baseline method and our proposed RAM method, respectively. Redder colors indicate greater influence.
@@
We also visualize the segmentation results of Nonlocal [29]/R50 [1] model with our proposed RAM and the baseline model under the attack setting of Permute. As shown in Fig. 4, most of the segmentation results towards the adversarial examples of our proposed RAM model still maintain the ground truth label. In contrast, the adversarial example almost completely modifies the output of the baseline model, and the segmentation results are very close to the target labels specified by the adversary. It demonstrates that the baseline model is quite vulnerable to patch-based attacks, while our RAM method significantly improves the robustness of semantic segmentation models against attacks. More visualization results are shown in Appendix B.
@@ A.1 The General Framework of Attack Methods
In the experiments, we set $T=400$ for all attack methods. For the hyperparameter of $\gamma$, we set $\gamma=0.005$ for PGD [13], and $\gamma=1$ for the other all attack methods.
@@ Appendix B More Visualization Results
[…]
(mIoU/%) (mIoU/%) (mIoU/%)
Baseline 48.25 49.32 40.01
RAD+MAS 48.16 45.61 35.17
MAS+RAD (ours) 48.30 33.72 22.63
Model Global Attention Local Attention
$\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
(pAcc/%) (pAcc/%) (pAcc/%) (pAcc/%) (pAcc/%) (pAcc/%) (pAcc/%) (pAcc/%)
Baseline 97.18 85.12 97.44 86.35 97.35 87.89 96.34 87.67
Certified [1*] 90.30 41.96 92.56 43.29 90.57 47.58 88.65 46.49
RAM (ours) 96.87 87.96 97.13 88.34 96.34 90.37 95.79 89.53